OpenStudy (anonymous):

A boat traveled 200 miles downstream and back. The trip downstream took 10 hours. The trip back took 70 hours. What is the speed of the boat in still water? What is the speed of the current?

hero (hero):

D = rt 200 = 10(r + c) 200 = 70(r - c) 200 = 10r + 10c 200 = 70r - 70c r = 20 - c r = 200/70 + c c = 60/7 r = 80/7 c = speed of current r = speed of boat c = 8.57 mph r = 11.43 mph

hero (hero):

I wonder if it was 100 miles there and 100 miles back. Doesn't seem clear

OpenStudy (vishal_kothari):

http://quizlet.com/4003656/math-practice-flash-cards/